Fermi National Laboratory
 

Operation Summary for October 27, 2010 - October 29, 2010

 

Wednesday Oct. 27

The day shift began with Operations monitoring the Tevatron (TeV) store 8210 (104.59E30), the Recycler stash (159.7E10), and the Antiproton Source (Pbar) stack (67.1E10), but with the Pbar target and PreVault areas in access for vacuum work, with continuing Pelletron work, with Linac sending beam to the Neutron Therapy Facility, with the Main Injector (MI) on, but not accelerating beam, with the Booster in access, with NuMI off for target work, with MiniBooNE in access, with the Meson FTBF CALICE T-978 experiment off due to no MI beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Booster came out of access at 10:50 AM. MiniBooNE came out of access at 11:10 AM. Operations resumed sending beam to MiniBooNE and T-978. Operations reported a problem with a Linac quadrupole power supply (QPS113). Operators searched and secured the PreVault area at 3:38 PM. Pbar resumed stacking at 3:50 PM.

During the evening shift, TeV personnel began Tevatron Electron Lens (TEL2) studies at 6:30 PM, which lasted until the end of the shift. TeV personnel also conducted crystal collimator studies. Operations disabled the T-978 beam at 6 PM. Operations reported that coalescing was unstable. Operators consulted MI and RF personnel; the problem lasted through to the end of shift. Operations slowed down the Pbar repetition rate due to the increasing amount of antiprotons stored, caused by the Pelletron work.

Thursday Oct. 28

The midnight shift began with Operations monitoring store 8210 (42.45E30) and the stash (156.8E10), with Pbar stacking (199.5E10), and with MiniBooNE taking beam.

During the midnight shift, Operations reported that coalescing was only marginal. TeV studies concluded at 1:25 AM. Operations terminated store 8210 at 1:35 AM and put the TeV into HEP shot setup at 1:45 AM, but Operations abandoned shot setup due to a faulty separator control card. Operators replaced the card, put the TeV into shot setup at 3:35 AM, and established store 8212 at 4:55 AM with an average initial luminosity (AIL) of 92E30. Operations reported that one transfer of antiprotons failed to coalesce and was not injected into the TeV, which made the store a 36x32 (36 bunches of protons by 32 bunches of antiprotons). Operations established beam to T-978 at 5:30 AM. Operations reported Booster RF problems.

During the day shift, Recycler personnel brought the Pelletron online at 9:25 AM; normal antiproton transfers to the Recycler resumed. At 9:50 AM, store 8212 aborted due to a separator spark. Operations reported that controls personnel traced the Booster RF sparking and to a faulty controls card. Also, MI personnel traced the previous shift's coalescing problems to a blown op-amp and a burned trace on the motherboard of an Internet Rack Monitor. Operations put the TeV into shot setup at 2:25 PM and established store 8213 at 3:55 PM with an AIL of 323E30. Operations disabled the T-978 beam when shot setup began.

During the evening shift, Safety System personnel re-cored the MI-65 doors to the interlocked key set. Operators searched and secured MI-65 and then conducted NuMI power supply testing.

Friday Oct. 29

The midnight shift began with Operations monitoring store 8213 (112.41E30) and the stash (228.1E10), with Pbar stacking (15.8E10), and with MiniBooNE taking beam.

During the midnight shift, Operations established beam to T-978 at 4:15 AM. Operations terminated store 8213 at 5 AM, put the TeV into shot setup, and established store 8214 at 6:15 AM with an AIL of 314E30. Operators prepared MI-65 for access at 6:30 AM.

The Plans for Friday and the Weekend

The plans for today are to stack, stash, and store. The Run Coordinator will allow MI personnel to conduct studies and for NuMI personnel to conduct target scans. NuMI plans to resume normal operations on Monday morning, 11/1/10.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/29/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory